Determine the value of i^10 as a real number.
Answer by stanbon(48545)

Determine the value of i^10 as a real number.
--------
i^4=1
i^2=-1
---------
i^10 = i^4 * i*4 * i^2 = 1*1*-1 = -1
============
Cheers,
Stan H.